# Simulation of charged nanotubes self-assembly during evaporation of a sessile droplet on a substrate

### v1

@@ Abstract
The ability to control the morphology of the nanotube deposit formed during the evaporation of a sessile droplet on a substrate is of theoretical and practical interest. Such deposits is required for various applications including nanotechnology, medicine, biotechnology, and optronics. In the experiment of Zhao et al. [J. Colloid Interface Sci. 440, 68 (2015)], an annular deposit was formed near the contact line. The deposition geometry is caused by the coffee-ring effect. This deposit is unusual in its morphology. It changes gradually in space from a disordered structure in the inner part of the ring to an aligned structure of nanotubes close to the periphery. To understand the mechanisms that lead to this, we have developed a mathematical model that takes into account the effects of advection, diffusion, and electrostatic interactions on particle transport. Results of numerical calculations have confirmed that all these factors together have an influence on the formation of such a variable morphology. Qualitative agreement with the experiment is shown for some values of the model parameters.
@@ I Introduction
Evaporation-induced self-assembly has promising for various applications in micro- and optoelectronics. For example, transparent electrically conductive films can be created from carbon nanotubes using inkjet printing [1]. This method is based on the coffee rings effect. In an experimental study [2] using this method, the possibility of creating a strain gauge was described. Different concentrations of surfactants in a solution containing carbon nanotubes were used and the concentration at which a wide annular deposit forms was determined. To increase the density of these deposits and improve electrical properties, droplets are applied alternately at the same location after the previous droplet has dried. This results in the formation of a dense multilayer ring deposit [2]. In another experiment [3], the effect of surfactant concentration and substrate material on the geometry of carbon nanotube deposits was studied. Different forms of deposition were obtained, including dot-like deposits, uniform deposits, coffee ring patterns, and combined patterns (coffee ring with a central dot-like deposit). The inkjet printing allows for the formation of relatively long and narrow layers of carbon nanotubes [4, 5]. The authors [4] studied the morphology of deposits formed near the contact line and measured their resistance. Thin-film transistors can be created based on these layers [5].
@@
In the experiment [6], carbon nanotubes were deposited during the movement of the contact line of a capillary bridge formed by the conclusion of a liquid layer between a substrate and the cover glass. The contact line moved in a <<stick-slip>> mode, leading to the formation of stripe-like deposits on the substrate. At the same time, there was periodic uneven detachment of the contact lines after the stage of pinning and spreading of its kink, with some particles being captured by the contact line during compression and others being transported by capillary flow caused by evaporation towards this boundary. Several parameters were studied: the height of the capillary bridge, the concentration of nanotubes and surfactants [6]. For coating and electronic component applications, it is sometimes useful to form a network of carbon nanotubes (uniformly distributed nanotubes) [7].The authors studied the effect of hydrophobic substrates and pre-centrifugation of the solution on the formation of deposits in the area of the dried droplet.
@@
The evaporative self-assembly of halloysite nanotubes was studied in a paper [8]. In their experiments, the authors observed the effect of coffee rings at the place where dried droplets were on the substrate. Key parameters such as the charge of the nanotubes, the temperature of the liquid, the aspect ratio of the nanotubes and the concentration of the solution were taken into account. At certain values of these parameters, a transition from chaos to ordering in the morphology of the annular deposit was occurred. Near the contact line, the orientation of the nanotubes was parallel to this boundary (nematic phase), and closer to the center there was a disordered deposition zone (isotropic phase). There was also an intermediate zone (transition phase). In another experiment with carbon nanotubes, the authors [9] studied the effect of particle concentration and substrate temperature on deposit morphology. Theoretically [10] and experimentally [11], it previously was shown that the transition from an isotropic to a nematic phase occurs for charged rod-like particles, including carbon nanotubes, occurs when the critical concentration of the solution is exceeded. The aim of this study was to numerically investigate the transition from a disordered to an aligned structures in order to gain a better understanding of the mechanisms that influence the morphology of nanotube deposits. The morphology of halloysite nanotube deposits is important, for example, in the development of membranes for dye/salt separation [12]. Halloysite clay tubes have been used as natural nanocontainers for loading and sustained release of chemical agents [13].
@@ II.1 Physical problem formulation
Consider a water droplet with suspended halloysite nanotubes evaporating on a solid horizontal impermeable substrate. The physical and geometric parameters of the problem are presented in the Table 1. The initial height of the drop $h_{0}$ is determined approximately from an experimental image (Fig. 2b from the Ref. [8]). Since the ratio between the height of the droplet and the radius of its base is small, $h_{0}/R\approx 0.3$, the droplet can be considered thin. The mass of a nanotube, $m$, was roughly estimated by multiplying its density by the volume of a cylinder-shaped tube. Throughout the process, the contact line remained pinned, as observed in Ref. [8], where the radius of the droplet base remained constant. In this case, non-uniform evaporation along the free surface of the droplet leads to capillary flow directed from its center to the periphery. This flow transfers the particles closer to the contact line, where they are gradually deposited onto the substrate (the coffee ring effect [14]). However, in this case, the situation is complicated by additional factors. Firstly, the particles are charged, which affects their dynamics. Secondly, they have an elongated shape. This particle shape affects their transport by flow and diffusion.
@@ II.2 Mathematical model
The process is described in a two-dimensional (2D) formulation, in the $xy$ plane, which is parallel to the substrate. Within the scope of this model, we will discuss the formation of a monolayer of nanotubes. However, it should be noted that multilayer deposits can also form in reality [15]. The effect of the transition from an isotropic to a nematic phase during evaporation of a droplet can be explained within the scope of such simple geometry. The review [16] on phase states of 2D elongated particle structures provides arguments in favor of the importance of studying such systems.
Within the scope of our model, we believe that nanotubes are characterized by a quasi-zero thickness. When checking for the absence of overlapping particles (collisions), we assume that the nanotubes have an infinite thickness in the horizontal plane. However, the particle thickness is taken into account in order to control the locations where the thickness of the liquid layer and the nanotube diameter are similar (vertical plane). This allows us to track the moment when a nanotube touches the substrate.
In this paper, a hybrid model is proposed for splitting by physical processes. Individual related processes can be approximated using different methods. Particle diffusion is modeled here by the Monte Carlo method. The direction of particle displacement and rotation at each time step is randomly chosen. The transfer of particles by flow, including their rotation, is described separately. Due to the low concentration and small size of the nanotubes, it is assumed that the particles have no effect on the fluid flow. When describing the electrostatic interaction between nanotubes within a single time step, we abstract from other transfer processes occurring simultaneously, but also take into account viscous friction. In the model, the different processes within a time step are executed sequentially. But at a sufficiently small time step ($\Delta t\to 0$), this model allows us to mimic the parallel processes in a real physical system.
@@ II.2.1 Droplet shape and fixing radius
Choose a formula to approximate the free surface of the droplet. Let’s check the parabolic approximation, formula (1) in Ref. [17], and the ellipsoidal approximation, formula (1) in Ref. [18]. We have determined the droplet shape using an experimental image (Fig. 2b in Ref. [8]) for the moment in time $t=$ 0 s and $t=$ 600 s. The radius of the droplet base (its value), $R$, is presented in Table 1. The results of the comparison are shown in Fig. 1. Both approximations provide similar values and are in good agreement with the experiment [8]. However, from a computational cost perspective, it is better to use the parabolic form [17]. Considering the parabolic shape, the expression for a fixing radius, $R_{f}$, was previously obtained [17]. It changes over time during evaporation of the droplet,
@@
where $\theta$ is the contact angle, $\theta(t)=\theta_{0}(1-t/t_{\mathrm{max}})$, $t_{\mathrm{max}}$ is the droplet evaporation time, and $d=l/\epsilon$ is the particle (cylinder) diameter. On this boundary, $R_{f}$, where the thickness of the liquid layer is commensurate with that of the nanotube, the particle is affected by the surface tension of the liquid. Thus, the nanotube is pressed against the substrate by a two-phase boundary (the <<liquid–air>> interface). In this model, it is assumed that the nanotube touching the substrate is fixed to a solid surface and remains stationary.
@@ II.2.2 Particle diffusion
When modeling the diffusion of nanotubes in a liquid, we assume that the solution is weak. Diffusion can be divided into two components: translational, with a diffusion coefficient $D_{t}=k_{B}T/f_{t}$ (m2/s), and rotational, with a diffusion coefficient $D_{r}=k_{B}T/f_{r}$ (1/s) [19]. In turn, translational diffusion is also conveniently divided into two components: along, $D_{\parallel}=k_{B}T/f_{\parallel}$, and perpendicular, $D_{\perp}=k_{B}T/f_{\perp}$, to the nanotube axis (Boltzmann constant $k_{B}\approx$ $1.38\times 10^{-23}$ J/K) [19]. Now let’s write the expressions for the Stokes friction coefficients [19],
$f_{t}=\frac{8\pi\eta l}{3\ln(\epsilon)+2\gamma_{\parallel}+\gamma_{\perp}},$
$f_{\parallel}=\frac{2\pi\eta l}{\ln(\epsilon)+\gamma_{\parallel}},$
$f_{\perp}=\frac{4\pi\eta l}{\ln(\epsilon)+\gamma_{\perp}},$
$f_{r}=\frac{\pi\eta l^{3}}{3(\ln(\epsilon)+\gamma_{r})}.$
@@ II.2.3 Effect of the fluid flow
Let us consider separately the possible types of motion of the nanotube under the influence of the fluid flow: (I) translational motion, (II) free rotational motion relative to the center of mass of the particle, and (III) rotational motion relative to a point of contact with another particle or the fixing radius, $R_{f}$. The radial flow velocity averaged over the thickness of the liquid layer will also be described, as it was done earlier [17],
@@
At each the $i$-th time step, the particle undergoes translational motion (I) with the flow velocity $\vec{v}$ at the point where the center (the point $\mathrm{C}$) of the nanotube mass is located. The velocity of the nanotube is recorded as a vector $\left.\vec{v}^{rod}=\vec{v}\left(\mathrm{C}\right)\right|_{i}$. Let’s write the velocity projections on the axis of the Cartesian coordinate system, $v_{x}=v_{r}\,x_{\mathrm{C}}/r$ and $v_{y}=v_{r}\,y_{\mathrm{C}}/r$. Over a period of time $\Delta t$ the nanotube moves by $\left.v^{rod}_{x}\right|_{i}\cdot\Delta t$ along the $OX$-axis and by $\left.v^{rod}_{y}\right|_{i}\cdot\Delta t$ along the $OY$-axis. Therefore, we use the following formulas in the calculation
$\left.x_{\mathrm{C}}\right|_{i+1}=\left.x_{\mathrm{C}}\right|_{i}+\Delta t\,\left.v^{rod}_{x}\right|_{i},\;\left.y_{\mathrm{C}}\right|_{i+1}=\left.y_{\mathrm{C}}\right|_{i}+\Delta t\,\left.v^{rod}_{y}\right|_{i}.$ (4)
The authors [9] hypothesized that the flow of liquid leads to the rotation of nanotubes during their convective transfer to the periphery of the droplet. According to this hypothesis, nanotubes should gradually align along the flow direction. Our numerical estimate has indicated that the torque is negligible, as the length of the nanotube is relatively short. Therefore, free rotational motion around the center of mass (II) can be ignored (see Appendix A).
If at some point in time one nanotube edge touches of another particle or fixing radius $R_{f}$, then further particle movement is only a rotational motion around the point of contact $A_{1}$ (III). The rotation angle of the nanotube during time $\Delta t$, around its edge (point $A_{1}$), on the current time step, is determined by formula (see Appendix A, formula (21))
$\left.\Delta\beta\right|_{i}\approx\left.\omega_{z}\right|_{i}\Delta t\approx\left.\frac{3\Delta t}{l^{2}}\left(\left(x_{C}-x_{A_{1}}\right)v_{y}(x_{C},\,y_{C})-\left(y_{C}-y_{A_{1}}\right)v_{x}(x_{C},\,y_{C})\right)\right|_{i}.$ (5)
The rotation is performed clockwise if $\Delta\beta>0$, and counterclockwise otherwise.
If a nanotube touches another particle, not with its edge, but at some point $N$, then the rotation of the nanotube in time $\Delta t$ around the point of contact at the calculated time step can be determined as (see Appendix A, formula (24))
$\left.\Delta\beta\right|_{i}\approx\frac{\Delta t}{l\cdot\left(\left|CN\right|^{2}+l^{2}/12\right)}\cdot\left(\left|A_{1}\,N\right|\cdot\left(\left(x_{Q}-x_{N}\right)v_{y}\left(x_{Q},\,y_{Q}\right)-\left(y_{Q}-y_{N}\right)v_{x}(x_{Q},\,y_{Q})\right)\right.+\\ +\left.\left.\left|A_{2}\,N\right|\cdot\left(\left(x_{P}-x_{N}\right)v_{y}(x_{P},\,y_{P})-\left(y_{P}-y_{N}\right)v_{x}(x_{P},\,y_{P})\right)\right)\right|_{i},$ (6)
where points $P$ and $Q$ are the midpoints of the segments into which the nanotube is divided by point $N$.

### v2

@@ Abstract
The ability to control the morphology of the nanotube deposit formed during the evaporation of a sessile droplet on a substrate is of theoretical and practical interest. Such deposits are required for various applications including nanotechnology, medicine, biotechnology, and optronics. In the experiment of Zhao et al. [J. Colloid Interface Sci. 440, 68 (2015)], an annular deposit was formed near the contact line. The deposition geometry is caused by the coffee ring effect. This deposit is unusual in its morphology. It changes gradually in space from a disordered structure in the inner part of the ring to an aligned structure of nanotubes close to the periphery. To understand the mechanisms that lead to this, we have developed a mathematical model that takes into account the effects of advection, diffusion, and electrostatic interactions on particle transport. Results of numerical calculations have confirmed that all these factors together have an influence on the formation of such a variable morphology. Qualitative agreement with the experiment is shown for some values of the model parameters.
@@ I Introduction
Evaporation-induced self-assembly has promising applications in micro- and optoelectronics. For example, transparent electrically conductive films can be created from carbon nanotubes using inkjet printing [1]. This method is based on the coffee ring effect. In an experimental study [2] using this method, the possibility of creating a strain gauge was described. Different concentrations of surfactants in a solution containing carbon nanotubes were used and the concentration at which a wide annular deposit forms was determined. To increase the density of these deposits and improve electrical properties, droplets are applied alternately at the same location after the previous droplet has dried. This results in the formation of a dense multilayer ring deposit [2]. In another experiment [3], the effect of surfactant concentration and substrate material on the geometry of carbon nanotube deposits was studied. Different forms of deposition were obtained, including dot-like deposits, uniform deposits, coffee ring patterns, and combined patterns (coffee ring with a central dot-like deposit). Inkjet printing enables the deposition of relatively long, narrow layers of carbon nanotubes [4, 5]. The authors [4] studied the morphology of deposits formed near the contact line and measured their resistance. Thin-film transistors can be created based on these layers [5].
@@
In the experiment with a capillary bridge formed by a confined liquid layer between the substrate and the cover glass [6], carbon nanotubes were deposited during the retraction of the contact line. The contact line moved in a stick-slip mode, leading to the formation of stripe-like deposits on the substrate. At the same time, there was periodic uneven detachment of the contact lines after the stage of pinning and spreading of its kink, with some particles being captured by the contact line during compression and others being transported by capillary flow caused by evaporation toward this boundary. Several parameters were studied: the height of the capillary bridge, the concentration of nanotubes and surfactants [6]. For coating and electronic component applications, it is sometimes useful to form a network of carbon nanotubes (uniformly distributed nanotubes) [7].The authors studied the effect of hydrophobic substrates and pre-centrifugation of the solution on the formation of deposits in the area of the dried droplet.
@@
The evaporative self-assembly of halloysite nanotubes was studied in a paper [8]. In their experiments, the authors observed the effect of coffee rings at the place where dried droplets were on the substrate. Key parameters such as the charge of the nanotubes, the temperature of the liquid, the aspect ratio of the nanotubes and the concentration of the solution were taken into account. At certain values of these parameters, a transition from chaos to ordering in the morphology of the annular deposit occurred. Near the contact line, the orientation of the nanotubes was parallel to this boundary (nematic phase), and closer to the center there was a disordered deposition zone (isotropic phase). There was also an intermediate zone (transition phase). In another experiment with carbon nanotubes, the authors [9] studied the effect of particle concentration and substrate temperature on deposit morphology. Theoretically [10] and experimentally [11], it was previously shown that the transition from an isotropic to a nematic phase occurs for charged rod-like particles, including carbon nanotubes, when the critical concentration of the solution is exceeded. The aim of this study is to numerically investigate the transition from a disordered to an aligned structure in order to gain a better understanding of the mechanisms that influence the morphology of nanotube deposits. The morphology of halloysite nanotube deposits is important, for example, in the development of membranes for dye/salt separation [12]. Halloysite clay tubes have been used as natural nanocontainers for loading and sustained release of chemical agents [13].
@@ II.1 Physical problem formulation
Consider a water droplet with suspended halloysite nanotubes evaporating on a solid horizontal impermeable substrate. The physical and geometric parameters of the problem are presented in Table 1. The initial height of the drop $h_{0}$ is determined approximately from an experimental image (Fig. 2(b) from Ref. [8]). Since the ratio between the height of the droplet and the radius of its base is small, $h_{0}/R\approx 0.3$, the droplet can be treated as a thin layer. The mass of a nanotube, $m$, was roughly estimated by multiplying its density by the volume of a cylinder-shaped tube. Throughout the process, the contact line remained pinned, as observed in Ref. [8], where the radius of the droplet base remained constant. In this case, non-uniform evaporation along the free surface of the droplet leads to capillary flow directed from its center to the periphery. This flow transfers the particles closer to the contact line, where they are gradually deposited onto the substrate (the coffee ring effect [14]). No Marangoni flow was detected in experiments using sessile water droplets with nearly identical geometry under comparable room conditions [15, 16, 8].
However, in this case, the situation is complicated by additional factors. First, the particles are charged, which affects their dynamics. Second, they have an elongated shape. This particle shape affects their transport by flow and diffusion. In the experiment [8], the nanotubes of halloysite clay after functionalization had a large surface charge and were repelled, i.e., their surface charge was not completely shielded in water.
@@ II.2 Mathematical model
The process is described using a two-dimensional (2D) formulation in the $xy$-plane, parallel to the substrate. Within the scope of this model, we will discuss the formation of a monolayer of nanotubes. However, it should be noted that multilayer deposits can also form in reality [17]. The effect of the transition from an isotropic to a nematic phase during the evaporation of a droplet can be explained within the scope of such simple geometry (see the sketch in Fig. 3 of Ref. [8]). The review [18] on phase states of 2D elongated particle structures provides arguments in favor of the importance of studying such systems.
Within the scope of our model, we assume that nanotubes are characterized by a quasi-zero thickness. When checking for the absence of particle overlap (collisions), we assume that the nanotubes are infinitely thin in the horizontal plane. However, the particle thickness is taken into account in order to control the locations where the thickness of the liquid layer and the nanotube diameter are similar (vertical plane). This allows us to track the moment when a nanotube touches the substrate. The approximation “zero-width objects” is often used for cases where $d\ll l$ (see, for example, the Ref. [19]).
In this paper, a hybrid model is proposed for splitting by physical processes. Individual related processes can be approximated using different methods. Particle diffusion is modeled here by the Monte Carlo method. The direction of particle displacement and rotation at each time step is randomly chosen. The transfer of particles by flow, including their rotation, is described separately. Due to the low concentration and small size of the nanotubes, it is assumed that the particles have no effect on the fluid flow (see Sec. 2 in Supplemental Material A). When describing the electrostatic interaction between nanotubes within a single time step, we abstract from other transfer processes occurring simultaneously, but also take into account viscous friction. In the model, the different processes within a time step are executed sequentially. But at a sufficiently small time step ($\Delta t\to 0$), this model allows us to mimic the parallel processes in a real physical system.
Let us consider the charge of nanotubes to be an effective value and a variable parameter. Here, we investigate the effect of this parameter on the ordering of the nanotubes near the droplet edge.
@@ II.2.1 Droplet shape and fixing radius
Choose a formula to approximate the free surface of the droplet. Let’s check the parabolic approximation, formula (1) in Ref. [20], and the ellipsoidal approximation, formula (1) in Ref. [21]. We have determined the droplet shape using an experimental image (Fig. 2( b) in Ref. [8]) for the moment in time $t=$ 0 s and $t=$ 600 s. The radius of the droplet base (its value), $R$, is presented in Table 1. A comparison of the results is shown in Fig. 1. Both approximations provide similar values and are in good agreement with the experiment [8]. However, the parabolic form is computationally more efficient [20]. Considering the parabolic shape, the expression for a fixing radius, $R_{f}$, was previously obtained [20]. It changes over time during evaporation of the droplet (see the sketch in Fig. 2 of Ref. [20]),
@@
where $\theta$ is the contact angle, $\theta(t)=\theta_{0}(1-t/t_{\mathrm{max}})$, $t_{\mathrm{max}}$ is the droplet evaporation time, and $d=l/\epsilon$ is the particle (cylinder) diameter. On this boundary, $R_{f}$, where the thickness of the liquid layer is commensurate with that of the nanotube, the particle is affected by the surface tension of the liquid. Thus, the nanotube is pressed against the substrate by a two-phase boundary (the “liquid–air” interface). This model assumes the nanotube in contact with the substrate remains fixed (stationary). It is worth noting that the formula (1) is limited by the condition $t\leqslant t_{\mathrm{cri}}=t_{\mathrm{max}}(1-2d/(\theta_{0}\,R))$. Otherwise we get a negative number under the root. In our case, the critical time is extremely close to the time of full evaporation, $t_{\mathrm{cri}}\approx 0.9999\,t_{\mathrm{max}}$. In addition, if the local height of a droplet is denoted as a function $h(r,t)$, then the nature of the boundary $R_{f}$ can be written mathematically as $h(R_{f}(t),t)=d$, where $r$ is the radial coordinate.
@@ II.2.2 Particle diffusion
When modeling nanotube diffusion in liquid, we assume a dilute solution. Diffusion can be divided into two components: translational, with a diffusion coefficient $D_{t}=k_{B}T/f_{t}$ (m2/s), and rotational, with a diffusion coefficient $D_{r}=k_{B}T/f_{r}$ (1/s) [22]. In turn, translational diffusion is also conveniently divided into two components: along, $D_{\parallel}=k_{B}T/f_{\parallel}$, and perpendicular, $D_{\perp}=k_{B}T/f_{\perp}$, to the nanotube axis (Boltzmann constant $k_{B}\approx$ $1.38\times 10^{-23}$ J/K) [22]. Now let’s write the expressions for the Stokes friction coefficients [22]:
$f_{t}=\frac{8\pi\eta_{0}l}{3\ln(\epsilon)+2\gamma_{\parallel}+\gamma_{\perp}},$
$f_{\parallel}=\frac{2\pi\eta_{0}l}{\ln(\epsilon)+\gamma_{\parallel}},$
$f_{\perp}=\frac{4\pi\eta_{0}l}{\ln(\epsilon)+\gamma_{\perp}},$
$f_{r}=\frac{\pi\eta_{0}l^{3}}{3(\ln(\epsilon)+\gamma_{r})}.$
@@ II.2.3 Effect of the fluid flow
Let us consider separately the possible types of motion of the nanotube under the influence of the fluid flow: (I) translational motion, (II) free rotational motion relative to the center of mass of the particle, and (III) rotational motion relative to a point of contact with another particle or the fixing radius, $R_{f}$. The radial flow velocity averaged over the thickness of the liquid layer will also be described, as was done earlier [20]:
@@
At each $i$-th time step, the particle undergoes translational motion (I) with the flow velocity $\vec{v}$ at the point where the center (point $\mathrm{C}$) of the nanotube mass is located. The velocity of the nanotube is recorded as a vector $\left.\vec{v}^{\mathrm{rod}}=\vec{v}\left(\mathrm{C}\right)\right|_{i}$. Let’s write the velocity projections on the axis of the Cartesian coordinate system, $v_{x}=v_{r}\,x_{\mathrm{C}}/r$ and $v_{y}=v_{r}\,y_{\mathrm{C}}/r$. Over a period of time $\Delta t$ the nanotube moves by $\left.v^{\mathrm{rod}}_{x}\right|_{i}\times\Delta t$ along the $OX$-axis and by $\left.v^{\mathrm{rod}}_{y}\right|_{i}\times\Delta t$ along the $OY$-axis. Therefore, we use the following formulas in the calculation
$\left.x_{\mathrm{C}}\right|_{i+1}=\left.x_{\mathrm{C}}\right|_{i}+\Delta t\,\left.v^{\mathrm{rod}}_{x}\right|_{i},\;\left.y_{\mathrm{C}}\right|_{i+1}=\left.y_{\mathrm{C}}\right|_{i}+\Delta t\,\left.v^{\mathrm{rod}}_{y}\right|_{i}.$ (4)
The authors [9] hypothesized that liquid flow induces nanotube rotation during convective transport toward the droplet periphery. According to this hypothesis, nanotubes should gradually align along the flow direction. Our numerical estimate has indicated that the torque is negligible, as the length of the nanotube is relatively short. Therefore, free rotational motion around the center of mass (II) can be ignored (see Appendix A).
When a nanotube edge contacts either another particle or the fixing radius $R_{f}$, its subsequent motion is constrained to rotation about the contact point $A_{1}$ (III). The rotation angle of the nanotube during time $\Delta t$, around its edge (point $A_{1}$), on the current time step, is determined by the formula [see Appendix A, formula (22)]
$\left.\Delta\beta\right|_{i}\approx\left.\omega_{z}\right|_{i}\Delta t\approx\\ \left.\frac{3\Delta t}{l^{2}}\left(\left(x_{C}-x_{A_{1}}\right)v_{y}(x_{C},\,y_{C})-\left(y_{C}-y_{A_{1}}\right)v_{x}(x_{C},\,y_{C})\right)\right|_{i}.$ (5)
The rotation is performed counterclockwise if $\Delta\beta>0$, and clockwise otherwise.
If a nanotube touches another particle, not with its edge, but at some point $N$, then the rotation of the nanotube in time $\Delta t$ around the point of contact at the calculated time step can be determined as [see Appendix A, formula (25)]
$\left.\Delta\beta\right|_{i}\approx\frac{\Delta t}{l\left(\left|CN\right|^{2}+l^{2}/12\right)}\times\\ \times\left(\left|A_{1}\,N\right|\times\left(\left(x_{Q}-x_{N}\right)v_{y}\left(x_{Q},\,y_{Q}\right)-\left(y_{Q}-y_{N}\right)v_{x}(x_{Q},\,y_{Q})\right)\right.+\\ +\left.\left.\left|A_{2}\,N\right|\times\left(\left(x_{P}-x_{N}\right)v_{y}(x_{P},\,y_{P})-\left(y_{P}-y_{N}\right)v_{x}(x_{P},\,y_{P})\right)\right)\right|_{i},$ (6)
where points $P$ and $Q$ are the midpoints of the segments into which the nanotube is divided by point $N$. Formula (5) is obtained from Formula (6) in the limit when $N\to A_{1}$.